A box contains 5 balls. Two are numbered 3, one is numbered 4, and two are numbered 5. The balls are mixed and one is selected 

A box contains 5 balls. Two are numbered 3, one is numbered 4, and two are numbered 5. The balls are mixed and one is selected 

at random. After a ball is selected, its number is recorded. Then it is replaced. If the experiment is repeated many times, 

at random. After a ball is selected, its number is recorded. Then it is replaced. If the experiment is repeated many times, 

find the mean of the numbers on the balls.


1
Expert's answer
2021-03-16T18:19:03-0400

Mean:

"\mu=\sum X\times P(X) = 3\times{\frac{2}{5}}+4\times{\frac{1}{5}}+5\times{\frac{2}{5}} = 4"

Answer: "\mu" = 4.